my father is 4 times older then me. in 20 years he will be only twice as old as me, how old am i and my father. what is the mathmatic problem and the steps to solve it
Found 2 solutions by TheProdicalSon, richwmiller:
Answer by TheProdicalSon(34)

set it up like this
4m = f
2m + 20 = f
Multiply the top equation by -1 to eliminate to f(represents father's age)
-4m = -f
2m + 20 = f
Add equations together
-2m + 20 = 0
-2m = -20
m = 10
You are 10 years old.
Your father is 40 years old.
Comment on my page if this was helpful

Answer by richwmiller(17219)

let f=father
let m=me
f=4m
f+20=2(m+20)
substitute f in second equation with 4m
4m+20=2(m+20)
4m+20=2m+40
subtract 2m from both sides
2m+20=40
subtract 20 from both sides
2m=20
divide by 2
m=10
f=4m
f=4*10=40
check
40+20=2(10+20)
40+20=2(30)=60
60=60
ok